OpenStudy (anonymous):

A force of 10N acts on a 5 kg mass for a short 0.5 s duration. What is the resulting velocity of the mass if it started with no velocity?

OpenStudy (anonymous):

Lets us consider Newton's Second Law, which mathematically boils down to \[F = m*a\] Where F is the force being applied to an object, m is the mass of the object and a is the acceleration of the object. The force (50 N) is being applied to a 5 kg object, so the resulting acceleration is: \[a = F/m = 10 m/s^2\] Should you integrate this with respect to time, you will find the velocity, due the the fact that acceleration is the derivative of velocity. So, using a simple integration, we find that:\[v = a*t +c\] Where a is the acceleration, t is the time, and c is a constant. Using the initial conditions, we find that v = 0 at t = 0. Using these values in the equation above gives c = 0. Therefore, the final velocity is: \[v = a*t = 10 m/s^2 * 0.5 s = 5 m/s\]

OpenStudy (anonymous):

Newton's Second Law is usually represented as, \[F=m*a\] but force can also be interpreted a change in momentum, mass*velocity.\[F = d(m*v)/dt\] By bringing the dt to the left hand side of the equation and integrating, one finds the new equation, \[F*t=m(v-v _{0})\] The quantity F*t is called impulse. The equation for impulse is more commonly written as \[I=m \Delta v\] Plugging in the values given \[10N*.5s=5kg(v-0)\] Therefor the final velocity is: \[v=1m/s\]